Ticket Number: HEA-NOC/20080408-3 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20080408 12:04 UTC+1 Problem Start: 20080408 13:00 UTC+1
Ticket Update: 20080408 22:05 UTC+1 Problem End: 20080408 20:41 UTC+1
Site/line: Blanch Services Network, Co-location, Web Hosting
Configuring BGP session to cr2-kp for the Services Network
Connectivity to the services network will be briefly affected while the
20080408 11:52 OMG The second 10Gbps link (mns4) from the Services network
in Blanchardstown to the new CRS in KP has been commissioned. BGP will be
configured to make this the primary link temporarily in advance of TG's
work tomorrow morning (ticket #20080327-6), as the current primary 1ink
(mns3) lands on the 10Gb card that is to be migrated to the second Bluenet
20080408 12:05 OMG Adding in colocation and webhosting customers to the
20080408 13:23 OMG BGP work on hold currently due to unrelated power
problems in the Blanchardstown PoP.
20080408 20:51 OMG BGP migration completed. Link to cr2-kp (mns4) is the
primary link, cr2-cwt (mns3) is the second and ar1-cwt (mns2) is the
tertiary link. Will leave ticket open for 24hours to monitor.
20080408 21:22 OMG Problems reaching some devices in KP from Blanch
(monitoring affected) and wit3-v6. Peerings shutdown while investigating
the problem. Suspect asymmetric routing.
20080408 21:56 OMG Will finish troubleshooting in the morning, the problem
appears to be related to 4 routes that are accepted on ar1-dcu, but not
on ar1-cwt, cr2-cwt and cr2-kp.
This ticket can be monitored at http://www.hea.net/tickets/20080408-3
HEAnet Limited. Registered in Ireland, No. 275301.